# Resilience for Regular Path Queries: Towards a Complexity Classification

### v2

@@ 3.1.2. Equivalence Between Local Languages and Letter-Cartesian Languages
By Proposition 3.6 we have that a language is local iff it is letter-Cartesian. So we show the equivalent claim: if $L$ is letter-Cartesian then so is $\mathtt{reduce}(L)$. Let $\alpha x\beta$ and $\gamma x\delta$ be in $\mathtt{reduce}(L)$, and let us show that $\alpha x\delta\in\mathtt{reduce}(L)$ as well. We have $\alpha x\beta\in L$ and $\gamma x\delta\in L$ and $L$ is letter-Cartesian, so we have $\alpha x\delta\in L$. Assume by contradiction that $\alpha x\delta\notin\mathtt{reduce}(L)$. Then $\alpha x\delta$ has a strict infix, call it $\tau$, that is in $L$. Notice that $\tau$ cannot be an infix of $\alpha$, because $\alpha x\beta\in\mathtt{reduce}(L)$ and $\tau$ would then be a strict infix of $\alpha x\beta$; and similarly $\tau$ cannot be an infix of $\delta$. Therefore $\tau$ must be of the form $\tau=\alpha_{1}x\delta_{1}\in L$ with $\alpha=\alpha_{2}\alpha_{1}$ and $\delta=\delta_{1}\delta_{2}$. We further have one of $\alpha_{2}\neq\epsilon$ and $\delta_{2}\neq\epsilon$, because $\tau$ is a strict infix of $\alpha x\delta$. Assume that $\alpha_{2}\neq\epsilon$ (the case of $\delta_{2}\neq\epsilon$ is symmetrical). Then we have $\alpha_{1}x\delta_{1}\in L$, and $\alpha x\beta\in L$, so since $L$ is letter-Cartesian we have $\alpha_{1}x\beta\in L$. But as $\alpha_{2}\neq\epsilon$ we know that $\alpha_{1}x\beta$ is a strict infix of $\alpha x\beta$, contradicting the fact that $\alpha x\beta\in\mathtt{reduce}(L)$. This concludes the proof. ∎
One related question is whether the criterion of Theorem 3.3 can easily be tested: can we easily test whether a given language $L$ is local? If $L$ is given as a DFA, then the task is in PTIME:
{toappendix}
3.3. Proof of Section 3.3: Locality Testing
{propositionrep}
[Locality testing] Given a DFA $A$, it is in PTIME to determine if $\mathcal{L}(A)$ is local.
@@
Given the DFA $A$, we build the RO-$\epsilon$NFA $A^{\prime}$ from Lemma 3.4. By this result and Proposition 3.6, we have $\mathcal{L}(A)\subseteq\mathcal{L}(A^{\prime})$, and $\mathcal{L}(A)$ is local iff $\mathcal{L}(A)=\mathcal{L}(A^{\prime})$. So to check whether $\mathcal{L}(A)$ is local, it is enough to check whether $\mathcal{L}(A^{\prime})\subseteq\mathcal{L}(A)$. For this, we complement $A$ in PTIME to build a DFA $A^{\prime\prime}$ with $\mathcal{L}(A^{\prime\prime})=\Sigma^{*}\setminus\mathcal{L}(A)$, and we must now check whether there is a word of $\mathcal{L}(A^{\prime})$ which is not in $\mathcal{L}(A)$, i.e., which is in $\mathcal{L}(A^{\prime\prime})$. So it suffices to check whether $\mathcal{L}(A)\cap\mathcal{L}(A^{\prime\prime})$ is empty or not. We can do this in PTIME using a product construction to build an $\epsilon$NFA recognizing $\mathcal{L}(A)\cap\mathcal{L}(A^{\prime\prime})$, and determining whether its language is empty or not by a reachability test. This concludes the proof. ∎
Note that Section 3.3 is about testing if the language $L=\mathcal{L}(A)$ recognized by $A$ is local (i.e., some local DFA for $L$ exists); it is not about testing if $A$ is a local DFA (which is easy in linear time).
If the input language $L$ is given as an NFA, we do not know the complexity of testing if $L$ is local: we show in Section 10 that it is PSPACE-hard to test if $L$ satisfies a property related to being local. Another question that remains open is to test whether the reduction of a given language $L$ is local: we can construct an automaton $A^{\prime}$ representing $\mathtt{reduce}(L)$ from an automaton $A$ representing $L$ and then study $A^{\prime}$, but the straightforward construction for $A^{\prime}$ entails an exponential blowup ((barcelo2019boundedness, Proposition 6)).
Read-Once automata.
Definition 3.7 (RO-$\epsilon$NFA).
A read-once $\epsilon$NFA (RO-$\epsilon$NFA for short) is an $\epsilon$NFA $A=(S,I,F,\Delta)$ such that for every letter $a\in\Sigma$ there is at most one transition of the form $(s,a,s^{\prime})$ in $\Delta$.
Example 3.8 (Example 3.2 continued).
The local DFA $A_{1}$ in Figure 2(a) for $ax^{*}b$ is an RO-$\epsilon$NFA, but the local DFA $A_{2}$ in Figure 2(b) for $ab|ad|cd$ is not. An RO-$\epsilon$NFA for this language is $A_{3}$ (Figure 2(c)).
We show that RO-$\epsilon$NFAs recognize precisely the local languages (and the presence of $\epsilon$-transitions is essential for this, see Section 11):
{toappendix}
3.4. Proof of Section 3.4: RO-$\epsilon$NFA = Local Languages
{lemmarep}
[RO-$\epsilon$NFA = local languages] A language $L$ is local iff it is recognized by an RO-$\epsilon$NFA. Moreover, given an $\epsilon$NFA $A$ recognizing a local language $L=\mathcal{L}(A)$, we can build in $O(|A|\times|\Sigma|)$ an RO-$\epsilon$NFA recognizing $L$.
Proof.
The first sentence follows from Section 3.1.1 and Section 3.1.1. The second sentence follows immediately from Proposition 3.6 and Lemma 3.4. ∎
The point of RO-$\epsilon$NFAs is that we can use them to solve resilience with a product construction:
@@
Given an $\epsilon$-NFA $A^{\prime}$ for the local language $L$, we compute an RO-$\epsilon$NFA $A=(S,I,F,\Delta)$ for $L$ in $O(|A^{\prime}|\times|\Sigma|)$ using Section 3.4. Let $D$ be the input bag database with fact multiplicities given by $\mathtt{mult}$, and let $V$ be the domain of $D$. Define a network $\mathcal{N}_{D,A}$ as follows:
@@
[…]
  • There is an edge with capacity +$\infty$+ ∞ from $(v,s)$ to $(v,s^{\prime})$ for all $v\in V$ and $s,s^{\prime}\in S$ such that $(s,\epsilon,s^{\prime})\in\Delta$.
  • There is an edge with capacity +$\infty$+ ∞ from $t_{\text{source}}$ to every vertex of the form $(v,s)$ with $s\in I$, and from every vertex of the form $(v,s)$ with $s\in F$ to $t_{\text{target}}$.
Clearly $\mathcal{N}_{D,A}$ is built in $O(|A|\times|D|)$ from $A$ and $D$. Moreover, there is a one-to-one correspondence between the contingency sets of $Q_{L}$ for $D$ and the cuts of $\mathcal{N}_{D,A}$ that have finite cost: intuitively, selecting facts of $D$ in the contingency set corresponds to cutting edges in $\mathcal{N}_{D,A}$, and ensuring that the remaining facts do not satisfy $Q_{L}$ corresponds to ensuring that no source to target paths remain. Crucially, as $A$ is read-once, the correspondence between facts of $D$ and edges of $\mathcal{N}_{D,A}$ with finite capacity is one-to-one. Further, the sum of multiplicities of the facts of a contingency set is equal to the cost of the corresponding cut, so minimum cuts correspond to minimum contingency sets. We can thus solve MinCut on $\mathcal{N}_{D,A}$ in $\tilde{O}(|\mathcal{N}_{D,A}|)$ (henzinger2024deterministic) to compute the answer to $\mathtt{RES}_{\mathrm{bag}}(L)$ in $D$. ∎
@@ 4. Hardness Techniques and First Results
In this section, we move on to cases where computing resilience is intractable. From now on we only study languages that are reduced, because showing hardness for a reduced language $L$ is sufficient to imply the same for any language $L^{\prime}$ such that $\mathtt{reduce}(L^{\prime})=L$.
@@
4.1. Hardness of $aa$
Let us first show that resilience is hard for the language $aa$. The result is known from the setting of CQs (DBLP:conf/pods/FreireGIM20, Prop. 10) by a direct reduction from 3-SAT. We give a self-contained proof where we reduce instead from minimum vertex cover: this proof serves as an intuitive presentation of the more general constructions that come next.
@@ 4.1. Hardness of $aa$
In other words, given a directed graph $G$ and integer $k\in\mathbb{N}$, it is NP-hard to decide if we can remove $k$ edges of $G$ such that $G$ no longer contains any path of length 2. We show Proposition 4.1 by reducing from the minimum vertex cover problem. Recall that a vertex cover of an undirected graph $G=(V,E)$ is a subset of vertices $V^{\prime}\subseteq V$ such that every edge of $E$ is incident to a vertex of $V^{\prime}$, i.e., for each $e\in E$, we have $e\cap V^{\prime}\neq\emptyset$. The vertex cover number of $G$ is the minimum cardinality of a vertex cover of $G$. Given an undirected graph $G$ and an integer $k\in\mathbb{N}$, the minimum vertex cover problem asks if the vertex cover number of $G$ is at most $k$. It is an NP-complete problem (karp1972reductibility).
(a) Completed gadget for $aa$
(b) Graph of $aa$-matches of Figure 3(a)
(c) Completed gadget for $axb|cxd$
(d) Condensed hypergraph of $axb|cxd$-matches of Figure 3(c)
Figure 3. Illustration of hardness gadgets and corresponding condensed hypergraphs of matches. Figures 3(a) and 3(b) illustrates Proposition 4.1, while Figures 3(c) and 3(d) illustrates Proposition 4.8.
@@
{toappendix}
4.2. Proof of Section 4.2: Hardness of Vertex Cover on Subdivisions
{propositionrep}
[Hardness of vertex cover on subdivisions] Let $\ell\in\mathbb{N}$ be an odd integer. Then for any undirected graph $G=(V,E)$, letting $G^{\prime}$ be an $\ell$-subdivision of $G$, then the vertex cover number of $G^{\prime}$ is $k+m(\ell-1)/2$, where $k$ is the vertex cover number of $G$ and $m$ is the number of edges of $G$.
@@
It suffices to show that $G$ has a vertex cover of cardinality at most $k$ iff $G^{\prime}$ has a vertex cover of cardinality at most $m(\ell-1)/2$.
We first show the first direction. Let $G=(V,E)$ and let $C$ be a vertex cover of $V$ of cardinality at most $k$. We pick a subset $C^{\prime}$ of the vertices of $G^{\prime}$ by taking the vertices of $C$ and picking vertices of the subdivisions in the following way. For each edge $\{u,v\}$ of $G$, letting $u-v_{1}-\cdots v_{\ell-1}-v$ be the corresponding path in $G$, we distinguish three cases:
@@
We next show the second direction, which is somewhat more challenging. Given a vertex cover $C^{\prime}$ of $G^{\prime}$ of cardinality at most $k+m(\ell-1)/2$, we will first modify $C^{\prime}$ without increasing its cardinality to guarantee the following property (*): for each edge $(u,v)\in E$, considering the corresponding path $u-v_{1}-\cdots v_{\ell-1}-v$ of $G^{\prime}$, then there are precisely $(\ell-1)/2$ of the intermediate vertices of the path that are in $C^{\prime}$, formally $C^{\prime}\cap\{v_{1},\ldots,v_{\ell-1}\}=(\ell-1)/2$.
For this, we process each path $u=v_{0}-v_{1}-\cdots v_{\ell-1}-v_{\ell}=v$, and consider each vertex $v_{1},\ldots,v_{\ell-1}$ in order. If for $i>1$ we have that $v_{i}\in C^{\prime}$ and the previous vertex $v_{i-1}$ was already in $C^{\prime}$, then we remove $v_{i}$ from $C^{\prime}$ and add instead $v_{i+1}$ to $C^{\prime}$. Note that the cardinality of $C^{\prime}$ does not increase, and may strictly decrease if $v_{i+1}$ was already in $C^{\prime}$. At the end of this process, given that we start with a vertex cover of $C^{\prime}$, we see that $C^{\prime}$ contains precisely $(\ell-1)/2$ vertices of the path: if $v_{1}\in C^{\prime}$ then we keep the $v_{i}$’s for which $i$ is odd, and otherwise the $v_{i}$’s for which $i$ is even.
We have processed $C^{\prime}$ to still be a vertex cover, and have cardinality $m(\ell-1)/2+k^{\prime}$ for some $k^{\prime}\leq k$, with $k^{\prime}=C^{\prime}\cap V$. Let us define $C\coloneq C^{\prime}\cap V$. It has cardinality at most $k$, and we claim that $C$ is a vertex cover. Indeed, assuming by contradiction that some edge $(u,v)$ of $G$ is not covered by $C$, so that $u\notin C$ and $v\notin C$, and considering the path $u=v_{0}-v_{1}-\cdots v_{\ell-1}-v_{\ell}=v$, we know that $C^{\prime}$ (after the processing of the previous paragraph) was a vertex cover of $G^{\prime}$, but if $u\notin C$ and $v\notin C$ then it cannot be the case that precisely $(\ell-1)/2$ of the $v_{i}$’s are in $C^{\prime}$ and cover all the edges of the path, a contradiction.
@@
(b) Database $\Xi$ constructed to encode $G^{\prime}$ with the gadget whose completion is pictured in Figure 3(a). All edges are labeled by the letter $a$. Letting $e=(u,v),e^{\prime}=(v,w)$ and $e^{\prime\prime}=(w,u)$, we write, e.g., $i^{\prime\prime}$ instead of $t_{e^{\prime\prime},i}$ for brevity.
Figure 4. Illustration of the reduction in the proof of Proposition 4.1.
[…]
The resilience of $aa$ on $\Xi$ is then equal to the vertex cover number of a 5-subdivision [^4] A simpler construction can achieve a 3-subdivision, but our construction will generalize better for subsequent definitions. of $G$. Indeed, define the graph of $aa$-matches on a database $D$ as the undirected graph having one vertex for each fact of $D$ and one edge for each pair of facts of $D$ that forms an $aa$-path. It is then clear that $\mathtt{RES}_{\mathrm{set}}(aa,\Xi)$ is equal to the vertex cover number of the graph of $aa$-matches of $\Xi$. Moreover, the graph of $aa$-matches of each $D_{e}^{\prime}$ is a path of length 5 from vertex $F_{e,\textup{in}}$ to vertex $F_{e,\textup{out}}$ (see Figure 3(b)). Thus, the graph of $aa$-matches of $\Xi$ is (isomorphic to) a 5-subdivision of $G$, and we conclude by Section 4.2. ∎

### v3

@@ 3.1. Local Languages
Let $L\coloneq\mathcal{L}(A)$. Build the local overapproximation $A^{\prime}$ of $L$ using Definition 3.8, which is clearly in PTIME. By 3.11, it suffices to check whether $L=\mathcal{L}(A^{\prime})$. It is well-known that testing equivalence of two DFAs is PTIME, so this concludes the proof. ∎
Note that Proposition 3.12 is about testing if the language $L=\mathcal{L}(A)$ recognized by $A$ is local (i.e., some local DFA for $L$ exists); it is not about testing if the specific DFA $A$ is a local DFA (which is easy to do in linear time simply by following the definition). It is also not about the complexity of testing whether the infix-free language of $\mathcal{L}(A)$ is local, which can be the case even if $\mathcal{L}(A)$ is not local. We can do this by building an automaton $A^{\prime}$ representing $\mathtt{IF}(L)$ from a DFA $A$ representing $L$ and then study $A^{\prime}$, but the straightforward construction for $A^{\prime}$ entails an exponential blowup ((barcelo2019boundedness, , Proposition 6)). We leave the precise complexity of this task open.
By contrast, if the input language $L$ is given as an NFA instead of a DFA, then testing if $L$ is local is PSPACE-complete (amarilli2025locality, ). This holds even under the assumption that $L$ is infix-free, so that the same lower bound applies to the problem of testing whether the infix-free language of an input NFA is local.
3.2. Tractability of Resilience for Local Languages
Having introduced local languages and presented their properties, we can now turn to the main result of this section, namely, that $\mathtt{RES}_{\mathrm{bag}}(L)$ is PTIME when $L$ is local:
Theorem 3.13 (Resilience for local languages).
[…]
Before proving Theorem 3.13, we remark that it implies a PTIME bound on $\mathtt{RES}_{\mathrm{bag}}(L)$ whenever $\mathtt{IF}(L)$ is local (as $\mathtt{RES}_{\mathrm{bag}}(L)$ is equivalent to $\mathtt{RES}_{\mathrm{bag}}(\mathtt{IF}(L))$). For instance, $\mathtt{RES}_{\mathrm{bag}}(L_{0})$ for $L_{0}=\{a,aa\}$ is in PTIME because $\mathtt{IF}(L_{0})=\{a\}$ is local, even though $L_{0}$ is not. In fact, to see if Theorem 3.13 applies to a language $L$, we can always focus on $\mathtt{IF}(L)$, thanks to the following lemma (whose proof is deferred to LABEL:apx:qwerty):
Lemma 3.14 (Infix-free sublanguages preserve locality).
If $L$ is local then so is $\mathtt{IF}(L)$.
[…]
Example 3.16 (Example 3.2 continued).
[…]
@@ 3.2. Tractability of Resilience for Local Languages
We first prove the first sentence of the claim.
[…]
The second sentence follows from the construction given in 3.7, noticing that the sets $\Sigma_{\mathrm{start}}$, $\Sigma_{\mathrm{end}}$ and $\Pi$ can be computed in PTIME from $A$. ∎
[…]
@@
[…]
@@
  • There is an edge with capacity $\mathtt{mult}(v\xrightarrow{a}v^{\prime})$ from $(v,s)$ to $(v^{\prime},s^{\prime})$ for all $v,v^{\prime}\in V$ and $s,s^{\prime}\in S$ and $a\in\Sigma$ such that $v\xrightarrow{a}v^{\prime}\in D$ and $(s,a,s^{\prime})\in\Delta$
  • There is an edge with capacity $\infty$ from $(v,s)$ to $(v,s^{\prime})$ for all $v\in V$ and $s,s^{\prime}\in S$ such that $(s,\epsilon,s^{\prime})\in\Delta$.
  • There is an edge with capacity $\infty$ from $t_{\text{source}}$ to every vertex of the form $(v,s)$ with $s\in I$, and from every vertex of the form $(v,s)$ with $s\in F$ to $t_{\text{target}}$.
Clearly $\mathcal{N}_{D,A}$ is built in $O(|A|\times|D|)$ from $A$ and $D$. Moreover, there is a one-to-one correspondence between the contingency sets of $D$ for $Q_{L}$ and the cuts of $\mathcal{N}_{D,A}$ that have finite cost: intuitively, selecting facts of $D$ in the contingency set corresponds to cutting edges in $\mathcal{N}_{D,A}$, and ensuring that the remaining facts do not satisfy $Q_{L}$ corresponds to ensuring that no source to target paths remain. Crucially, as $A$ is read-once, the correspondence between facts of $D$ and edges of $\mathcal{N}_{D,A}$ with finite capacity is one-to-one. Further, the sum of multiplicities of the facts of a contingency set is equal to the cost of the corresponding cut, so minimum cuts correspond to minimum contingency sets. We can thus solve MinCut on $\mathcal{N}_{D,A}$ in $\tilde{O}(|\mathcal{N}_{D,A}|)$ (henzinger2024deterministic, ) to compute the answer to $\mathtt{RES}_{\mathrm{bag}}(L)$ in $D$. ∎
@@ 4. Hardness Techniques and First Results
In this section, we move on to cases where computing resilience is intractable. From now on we only study languages that are infix-free, because showing hardness for an infix-free language $L$ is sufficient to imply the same for any language $L^{\prime}$ such that $\mathtt{IF}(L^{\prime})=L$.
@@
4.1. Hardness of aaaa
Let us first show that resilience is hard for the language $aa$. The result is known from the setting of CQs (DBLP:conf/pods/FreireGIM20, , Prop. 10) by a direct reduction from 3-SAT. We give a self-contained proof where we reduce instead from minimum vertex cover: this proof serves as an intuitive presentation of the more general constructions that come next.
@@ 4.1. Hardness of aaaa
In other words, given a directed graph $G$ and integer $k\in\mathbb{N}$, it is NP-hard to decide if we can remove $k$ edges of $G$ such that $G$ no longer contains any walk of length 2. We show Proposition 4.1 by reducing from the minimum vertex cover problem. Recall that a vertex cover of an undirected graph $G=(V,E)$ is a subset of vertices $V^{\prime}\subseteq V$ such that every edge of $E$ is incident to a vertex of $V^{\prime}$, i.e., for each $e\in E$, we have $e\cap V^{\prime}\neq\emptyset$. The vertex cover number of $G$ is the minimum cardinality of a vertex cover of $G$. Given an undirected graph $G$ and an integer $k\in\mathbb{N}$, the minimum vertex cover problem asks if the vertex cover number of $G$ is at most $k$. It is an NP-complete problem (karp1972reductibility, ).
@@
Proposition 4.2 (Hardness of vertex cover on subdivisions).
Let $\ell\in\mathbb{N}$ be an odd integer. Then for any undirected graph $G=(V,E)$, letting $G^{\prime}$ be an $\ell$-subdivision of $G$, then the vertex cover number of $G^{\prime}$ is $k+m(\ell-1)/2$, where $k$ is the vertex cover number of $G$ and $m$ is the number of edges of $G$.
@@
It suffices to show that $G$ has a vertex cover of cardinality at most $k$ if and only if $G^{\prime}$ has a vertex cover of cardinality at most $k+m(\ell-1)/2$.
We first show the “only if” direction. Let $G=(V,E)$ and let $C$ be a vertex cover of $V$ of cardinality at most $k$. We pick a subset $C^{\prime}$ of the vertices of $G^{\prime}$ by taking the vertices of $C$ and picking vertices of the subdivisions in the following way. For each edge $\{u,v\}$ of $G$, letting $u-v_{1}-\cdots v_{\ell-1}-v$ be the corresponding path in $G$, we distinguish three cases:
@@
We next show the second direction, which is somewhat more challenging. Given a vertex cover $C^{\prime}$ of $G^{\prime}$ of cardinality at most $k+m(\ell-1)/2$, we will first modify $C^{\prime}$ without increasing its cardinality to guarantee the following property (*): for each edge $(u,v)\in E$, considering the corresponding path $u-v_{1}-\cdots v_{\ell-1}-v$ of $G^{\prime}$, then there are precisely $(\ell-1)/2$ of the intermediate vertices of the path that are in $C^{\prime}$, formally $|C^{\prime}\cap\{v_{1},\ldots,v_{\ell-1}\}|=(\ell-1)/2$.
For this, we process each path $u=v_{0}-v_{1}-\cdots v_{\ell-1}-v_{\ell}=v$, and consider each vertex $v_{1},\ldots,v_{\ell-1}$ in order. If for $i>1$ we have that $v_{i}\in C^{\prime}$ and the previous vertex $v_{i-1}$ was already in $C^{\prime}$, then we remove $v_{i}$ from $C^{\prime}$ and add instead $v_{i+1}$ to $C^{\prime}$. Note that the cardinality of $C^{\prime}$ does not increase, and may strictly decrease if $v_{i+1}$ was already in $C^{\prime}$. At the end of this process, given that we start with a vertex cover $C^{\prime}$ of $G^{\prime}$, we see that $C^{\prime}$ contains precisely $(\ell-1)/2$ vertices of the path: if $v_{1}\in C^{\prime}$ then we keep the $v_{i}$’s for which $i$ is odd, and otherwise the $v_{i}$’s for which $i$ is even.
We have processed $C^{\prime}$ to still be a vertex cover, and have cardinality $m(\ell-1)/2+k^{\prime}$ for some $k^{\prime}\leq k$, with $k^{\prime}=|C^{\prime}\cap V|$. Let us define $C\coloneq C^{\prime}\cap V$. It has cardinality at most $k$, and we claim that $C$ is a vertex cover. Indeed, assuming by contradiction that some edge $(u,v)$ of $G$ is not covered by $C$, so that $u\notin C$ and $v\notin C$, and considering the path $u=v_{0}-v_{1}-\cdots v_{\ell-1}-v_{\ell}=v$, we know that $C^{\prime}$ (after the processing of the previous paragraph) was a vertex cover of $G^{\prime}$, but if $u\notin C$ and $v\notin C$ then it cannot be the case that precisely $(\ell-1)/2$ of the $v_{i}$’s are in $C^{\prime}$ and cover all the edges of the path, a contradiction.
@@
(c) Graph of $aa$-matches of Figure 3(b)
(d) Database $\Xi$ constructed to encode $G^{\prime}$. All edges are labeled by $a$.
Figure 3. Illustration of the proof of Proposition 4.1.
(b) Completed gadget for aaaa
(a) Completed gadget for $axb|cxd$
(b) Hypergraph of matches of $axb|cxd$ on the completed gadget in Figure 4(a)
(c) Condensed hypergraph of matches of Figure 4(b)
Figure 4. Illustration of hardness gadgets and corresponding condensed hypergraphs of matches for the proof of Proposition 4.13.
We reduce from the vertex cover problem. Let $G$ be an undirected graph and $k\in\mathbb{N}$ be an integer. We perform the following construction, illustrated in Figure 3(a). We start by picking an arbitrary orientation for each edge of $G$, giving a directed graph $G^{\prime}=(V,E)$: see Figure 3(a). We then construct a database $\Xi$ in the following way. First, for each vertex $u\in V$ we create a fact $s_{u}\xrightarrow{a}t_{u}$. Second, for each edge $e=(u,v)$ of $E$ we create the following set $D_{e}$ of facts: $t_{u}\xrightarrow{a}t_{e,1}$, $t_{e,1}\xrightarrow{a}t_{e,2}$, $t_{e,2}\xrightarrow{a}t_{e,3}$, $t_{v}\xrightarrow{a}t_{e,2}$. Together with $s_{u}\xrightarrow{a}t_{u}$ and $s_{v}\xrightarrow{a}t_{v}$, the facts of $D_{e}$ form an instance that we call the completed gadget $D_{e}^{\prime}$ for $e$: we illustrate $D_{e}^{\prime}$ in Figure 3(b), abbreviating $t_{e,i}$ as $i$ for brevity. We introduce the notations $F_{e,\textup{in}}=s_{u}\xrightarrow{a}t_{u}$ and $F_{e,\textup{out}}=s_{v}\xrightarrow{a}t_{v}$ to refer to the two facts of $D_{e}^{\prime}\setminus D_{e}$, called the endpoint facts: they are shown in bold orange in Figure 3(b). The overall database $\Xi$ built from the orientation $G^{\prime}$ of $G$ is illustrated in Figure 3(d), where for brevity we abbreviate $t_{e,i}$ as $i_{e}$. Building $\Xi$ from $G$ is in $O(|G|)$ (as $D_{e}$ has constant size).
We then claim that the resilience of $aa$ on $\Xi$ is equal to the vertex cover number of a 5-subdivision of $G$. Indeed, define the graph of $aa$-matches on a database $D$ as the undirected graph having one vertex for each fact of $D$ and one edge for each pair of facts of $D$ that forms an $aa$-walk. It is then clear that $\mathtt{RES}_{\mathrm{set}}(aa,\Xi)$ is equal to the vertex cover number of the graph of $aa$-matches of $\Xi$. Moreover, the graph of $aa$-matches of each $D_{e}^{\prime}$ is a path of length 5 from vertex $F_{e,\textup{in}}$ to vertex $F_{e,\textup{out}}$ (see Figure 3(c)). Thus, the graph of $aa$-matches of $\Xi$ is (isomorphic to) a 5-subdivision of $G$, and we conclude by Proposition 4.2. ∎